If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
The answer is 1.09 days, however I need to figure out the mathmatical breakdown that gives me this answer as my equations give me 1.23 days.
4/6=0.7
6/2=3
3.7/3=1.23
Any help would be greatly appreciated!!
Thank you,
Erica

Answer by Alan3354(69443)

If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
------------------
You have to use the reciprocals:
Sam does 1/4 of the job per day
Lisa does 1/6
Tom does 1/2
Add those up.
1/4 + 1/6 + 1/2 = 11/12 job/day together.
Then, it takes 12/11 days to do the job.
12/11 = 1.090909... days.